what is the length of the transverse axis of the hyperbola whose equation is (x-7)^2/9^2-(y-4)^2/5^2=1
Answer by lwsshak3(11628)

what is the length of the transverse axis of the hyperbola whose equation is
(x-7)^2/9^2-(y-4)^2/5^2=1
This is an equation for a hyperbola with horizontal transverse axis of the standard form:
(x-h)^2/a^2-(y-k)^2/b^2=1, (h,k) being the (x,y) coordinates of the center.
..
a^2=9^2
a=9
length of horizontal transverse axis=2a=18